# Breaking the Performance Ceiling in Complex Reinforcement Learning requires Inference Strategies

### v1

@@ Acknowledgements
We would like to thank Guillaume Toujas-Bernate, Jake Lourie and Thomas Lecat for useful discussions on the use of inference strategies in real-world applications. We thank our MLOps team for developing our model training and experiment orchestration platform AIchor. We thank the Python and JAX communities for developing tools that made this research possible. Finally, we thank Google’s TPU Research Cloud (TRC) for supporting our research with Cloud TPUs.
@@ Appendix A Additional results
Section 4.2 presents the contour plots of Sable with all inference strategies on the hardest 8 tasks of the benchmark. We report the contour plots of the other $9$ remaining tasks on Fig. 8.
@@
Fig. 8, shows similar trends to those illustrated in Section 4.2. sable +compass leads to the best overall performance. Stochastic sampling provides a good and robust baseline overall. Finally, sgbs and online fine-tuning suffer from variance: despite being able to compete close to compass on a couple tasks, they often get outperformed by stochastic sampling. We do not report the results for sable +sgbs on smacv2_10_units because we found an issue, but the figure will be updated shortly with new results.
@@ Appendix B Details about the tasks
Figure 9: Environment rendering for Connector. Task name: con-10x10-10a. Image from Mahjoub et al. [2025].
@@
is an environment designed to model the routing of a printed circuit board (PCB). Agents start at randomly generated initial positions and have to reach a goal position without overlapping with each other’s paths (indicated as lower opacity coloured cells in Figure 9). Each agent has a partial view with a fixed field of view around itself as well as its current $(x,y)$-coordinate on the grid and the $(x,y)$-coordinate of its goal location. At each timestep, agents receive a small negative reward $-0.03$ and receive a reward of $1$ for successfully connecting to a goal location. The particular difficulty in this environment stems from the fact that agents have to select actions carefully so that they not only reach their goal greedily but so that all agents can ultimately reach their goals.
Figure 10: Environment rendering for Robot Warehouse. Task name: tiny-2ag. Image from Mahjoub et al. [2025].
@@
is an environment where agents need to coordinate to deliver shelves (green cells in Figure 10) to a goal location (dark grey cells in Figure 10) in a warehouse. The reward is sparse since agents only receiving a reward of $1$ for a successful shelf delivery to the goal location and 00 otherwise. This sparsity makes the environment particularly difficult since agents have to complete a long sequence of correct actions in order to receive any reward signal.
Figure 11: Environment rendering for SMAX. Task name: 2s3z. Image from Mahjoub et al. [2025].
@@
is an environment where agents need to coordinate and develop the correct micro strategies in order to defeat an enemy team. It was shown that the original benchmark was overly simplistic and for this reason SMACv2 [Ellis et al., 2023] was developed in order to address the shortcomings of the original benchmark. SMACv2 randomly generates a team of allied units by sampling from a selection of unit types and randomly generates ally starting positions at the start of each episode. The need for generalisation is what gives this benchmark its difficulty.
@@
Since Connector was developed as part of Jumanji [Bonnet et al., 2023], we make use of the JAX-based implementation directly. For RWARE we use the JAX-based reimplementation from Jumanji and for SMAC and SMACv2 we make use of the JAX-based reimplementation from JaxMARL [Rutherford et al., 2023], named SMAX. For a detailed discussion on task naming conventions we refer the reader to the Appendix of Mahjoub et al. [2025].
Appendix C Hyper-parameters
@@
Hyper-parameters used to train the base policies for this benchmark are taken from Mahjoub et al. [2025]. These hyper-parameters were tuned on each task with a budget of 40 trials using the Tree-structured Parzen Estimator (TPE) Bayesian optimisation algorithm, and are reported in Mahjoub et al. [2025].
@@
Hyper-parameters used to train all the compass checkpoints can be found in Table 1. Most of them were kept close to the original hyper-parameters used in Chalumeau et al. [2023b].
@@
Most values are directly taken from original works or kept close to these. When different, we tried to use guidance from original work to decide how to set them. The values for the cma-es search used with compass, for all base policies, can be found in Table 2. The values for online fine-tuning (all base policies) are reported in Table 3. Hyper-parameters for sgbs (all base policies) are reported in Table 4, and for stochastic sampling in Table 5.
@@
Table 4: Hyper-parameters for Simulation Guided Beam Search (sgbs). The same values are used for sable, ippo and mappo. Values are chosen such that beam width and top-k use fully the number of attempts allowed by batch.
@@
Beam width 4 4 4 4 4 4
Top k 1 4 8 16 32 64
@@
Appendix D Experimental stack
All algorithmic implementations are built by extending the JAX-based research library, Mava [de Kock et al., 2021]. Our version of cma-es used for searching compass’s latent space comes from the JAX-based quality diversity library, QDAX [Chalumeau et al., 2024a]. Both these libraries are built to leverage the DeepMind JAX ecosystem [DeepMind et al., 2020] and use Flax [Heek et al., 2024] for building neural networks, optax for optimisers and orbax for model checkpointing.
Appendix E Additional details about training
E.1 Base policy training
In Section 3, we explain that in our setting, it is assumed that all methods can be trained until convergence. For transparency and comparison with previous results in the literature, we report in  Table 6 the training budget given to the methods to reach convergence, in environment steps.
@@ E.1 Base policy training
E.2 COMPASS training
In this section, we provide details about compass and our implementation. For a more in-depth explanation, we refer the reader to the original paper [Chalumeau et al., 2023b]. Our implementation is available in the submitted code files.
@@ E.2 COMPASS training
In the original work, the additional weights added to the base neural network are initialised with zeros. In our case, we observed that this could not provide enough diversification in the first training step, preventing any emergence of specialisation, leading to no benefit. We found that initialising these parameters with random uniform values between $-0.01$ and $0.01$ fixed the issue across all tasks.
@@
Similarly to the original method, the latent space is not learned, it is a fixed prior, and the network’s weights are learned to create diversity from this prior space. We also use a similar prior, which is a uniform distribution between $-1$ and $1$, over $16$ dimensions. We do not need any amplification of the latent vector (i.e., we multiply by $1$, whereas the original work multiplies it by $100$). At each training step, a batch of instances is sampled from the training distribution, a batch of latent vectors is sampled from that latent space, the conditioned policy is evaluated for each instance, for each latent vector, and only the best performing latent vectors, for each instance, are used in the computation of the loss. This creates the specialisation (diversification). For additional motivation and mathematical formulation, we refer the reader to Chalumeau et al. [2023b]. Like all training processes in our experimental study, the method is allowed to train until convergence. In practice, each compass checkpoint was trained for $100$ million steps.
Appendix F Pseudo-algorithms
In this section, we provide algorithmic descriptions for the inference strategies used in the paper, showing how the policy is used at inference time, under the time and budget constraints. Stochastic sampling is explained in Algorithm 1, sgbs in Algorithm 2, online fine-tuning in Algorithm 3 and compass in Algorithm 4.
@@ Appendix F Pseudo-algorithms
3:  while elapsed time $<T_{\max}$ do
@@
3:  while elapsed time $<T_{\max}$ do
@@
4:  while elapsed time $<T_{\max}$ do
@@
4:  while elapsed time $<T_{\max}$ do
@@
Appendix G Discussion about the inference-time budget
@@
To remain consistent with previous literature, and provide a better understanding of the search and adaptation cost of each inference strategy and each base policy, we provide the number of attempts achieved within the time limit, for all tasks and compute capacity (i.e., number of parallel attempts allowed). Values for sable within 300 seconds are reported on Table 7 for compass, Table 8 for online fine-tuning, Table 9 for stochastic sampling. sgbs builds new solutions from existing partial solutions, hence the count of the attempted solution cannot be obtained in the same way as other methods.
Values for all base policies within 30 seconds are reported on Table 10Table 12 and Table 11.

### v3

@@ Acknowledgements
We would like to thank Guillaume Toujas-Bernate, Jake Lourie and Thomas Lecat for useful discussions on the use of inference strategies in real-world applications. We thank our MLOps team for developing our model training and experiment orchestration platform AIchor. We thank the Python and JAX communities for developing tools that made this research possible. We thank the anonymous reviewers for their constructive feedback and valuable suggestions. Finally, we thank Google’s TPU Research Cloud (TRC) for supporting our research with Cloud TPUs.
@@ Appendix A Additional results
Section˜4.2 presents the contour plots of Sable with all inference strategies on the hardest 8 tasks of the benchmark. We report the contour plots of the other $9$ remaining tasks on Fig.˜8.
@@
Fig.˜8, shows similar trends to those illustrated in Section˜4.2. sable +compass leads to the best overall performance. Stochastic sampling provides a good and robust baseline overall. Online fine-tuning suffer from variance: despite being able to compete close to compass on a couple tasks, it often get outperformed by stochastic sampling. Finally, sgbs obtains good performance on numerous scenarios, and even leads one scenario (smacv2_10_units).
@@ Appendix B Details about the tasks
Figure 9: Environment rendering for Connector. Task name: con-10x10-10a. Image from mahjoub2025sable.
@@
is an environment designed to model the routing of a printed circuit board (PCB). Agents start at randomly generated initial positions and have to reach a goal position without overlapping with each other’s paths (indicated as lower opacity coloured cells in Figure 9). Each agent has a partial view with a fixed field of view around itself as well as its current $(x,y)$-coordinate on the grid and the $(x,y)$-coordinate of its goal location. At each timestep, agents receive a small negative reward 0.03-0.03 and receive a reward of $1$ for successfully connecting to a goal location. The particular difficulty in this environment stems from the fact that agents have to select actions carefully so that they not only reach their goal greedily but so that all agents can ultimately reach their goals.
Figure 10: Environment rendering for Robot Warehouse. Task name: tiny-2ag. Image from mahjoub2025sable.
@@
is an environment where agents need to coordinate to deliver shelves (green cells in Figure 10) to a goal location (dark grey cells in Figure 10) in a warehouse. The reward is sparse since agents only receiving a reward of $1$ for a successful shelf delivery to the goal location and 0 otherwise. This sparsity makes the environment particularly difficult since agents have to complete a long sequence of correct actions in order to receive any reward signal.
Figure 11: Environment rendering for SMAX. Task name: 2s3z. Image from mahjoub2025sable.
@@
is an environment where agents need to coordinate and develop the correct micro strategies in order to defeat an enemy team. It was shown that the original benchmark was overly simplistic and for this reason SMACv2 [ellis2023smacv2] was developed in order to address the shortcomings of the original benchmark. SMACv2 randomly generates a team of allied units by sampling from a selection of unit types and randomly generates ally starting positions at the start of each episode. The need for generalisation is what gives this benchmark its difficulty.
@@
Since Connector was developed as part of Jumanji [jumanji2023github], we make use of the JAX-based implementation directly. For RWARE we use the JAX-based reimplementation from Jumanji and for SMAC and SMACv2 we make use of the JAX-based reimplementation from JaxMARL [rutherford2023jaxmarl], named SMAX. For a detailed discussion on task naming conventions we refer the reader to the Appendix of mahjoub2025sable.
Appendix C Results and discussion about the single-agent case
C.1 Additional results on a single-agent challenge: Craftax
To demonstrate that our core message is not limited to the multi-agent case, we selected a single-agent task that is still far from being solved by RL algorithms, Craftax [matthews2024craftax].
We trained a sable base policy and matched the current state-of-the-art performance reported on the Craftax-1B leaderboard, which is 18.318.3% of the maximum score. We then ran a $30$-second inference-time search with the simplest inference strategy (stochastic sampling). The results are reported in Table˜1. Strikingly, and similar to our main results, using only a $30$-second inference-time search yields a 3737% performance gain, reaching 25.825.8% of the maximum score. To further put this into context, this performance boost is of the same magnitude as that of the last $300$ million steps of training the base policy.
These new results confirm that the significant improvement enabled by inference-time search is not limited to multi-agent RL and is applicable to RL in general. In the next subsection, we discuss existing results (reported in the literature) to further support the importance of inference-time search in single-agent settings.
Table 1: Performance comparison on Craftax-1B between zero-shot inference and stochastic sampling with a 30-second budget. Values indicate mean performance with 95% confidence intervals.
Scenario Zero-shot Stochastic Sampling (30s budget)
Craftax-1B 18.8 (18.1, 19.5) 25.8 (25.7, 26.1)
C.2 Published results on single-agent tasks
To further demonstrate the impact of inference-time strategies beyond the multi-agent case, we discuss previous results reported in the literature. These results do not show the same impact because they are often close to saturation, even with zero-shot. They nevertheless demonstrate that inference-time strategies can be successfully applied to a wide range of problems and provide noticeable improvement.
We refer to the experimental results reported in the seminal works that introduced the inference strategies used in our paper [chalumeau2023combinatorial, choo2022simulationguided, Bello16]. All of these evaluate on single-agent tasks (TSP, CVRP, JSSP, Knapsack, FFSP) and show that inference strategies bring statistically significant improvements over zero-shot performance.
However, we recall the limitations we have identified on these experimental results: the limited breadth of the time/compute settings, methods’ reliance on problem-specific tricks and saturated benchmarks where zero-shot performance is already very high.
Appendix D Hyperparameters
@@
Hyper-parameters used to train the base policies for this benchmark are taken from mahjoub2025sable. These hyper-parameters were tuned on each task with a budget of 40 trials using the Tree-structured Parzen Estimator (TPE) Bayesian optimisation algorithm, and are reported in mahjoub2025sable.
@@
Hyper-parameters used to train all the compass checkpoints can be found in Table˜2. Most of them were kept close to the original hyper-parameters used in chalumeau2023combinatorial.
@@
Most values are directly taken from original works or kept close to these. When different, we tried to use guidance from original work to decide how to set them. The values for the cma-es search used with compass, for all base policies, can be found in Table˜3. The values for online fine-tuning (all base policies) are reported in Table˜4. Hyper-parameters for sgbs (all base policies) are reported in Table˜5, and for stochastic sampling in Table˜6.
@@
Table 5: Hyper-parameters for Simulation Guided Beam Search (sgbs). The same values are used for sable, ippo and mappo. Values are chosen such that expansion factor and beam width (top-k) use fully the number of attempts allowed by batch.
@@
Expansion factor 4 4 4 4 4 4
Beam width (Top k) 1 4 8 16 32 64
@@
Appendix E Experimental stack
All algorithmic implementations are built by extending the JAX-based research library, Mava [de2021mava]. Our version of cma-es used for searching compass’s latent space comes from the JAX-based quality diversity library, QDAX [chalumeau2024qdax]. Both these libraries are built to leverage the DeepMind JAX ecosystem [deepmind2020jax] and use Flax [flax2020github] for building neural networks, optax for optimisers and orbax for model checkpointing.
Appendix F Additional details about training
F.1 Base policy training
In Section˜3, we explain that in our setting, it is assumed that all methods can be trained until convergence. For transparency and comparison with previous results in the literature, we report in  Table˜7 the training budget given to the methods to reach convergence, in environment steps.
@@ F.1 Base policy training
F.2 COMPASS training
In this section, we provide details about compass and our implementation. For a more in-depth explanation, we refer the reader to the original paper [chalumeau2023combinatorial]. Our implementation is available in the submitted code files.
@@ F.2 COMPASS training
In the original work, the additional weights added to the base neural network are initialised with zeros. In our case, we observed that this could not provide enough diversification in the first training step, preventing any emergence of specialisation, leading to no benefit. We found that initialising these parameters with random uniform values between 0.01-0.01 and $0.01$ fixed the issue across all tasks.
@@
Similarly to the original method, the latent space is not learned, it is a fixed prior, and the network’s weights are learned to create diversity from this prior space. We also use a similar prior, which is a uniform distribution between 1-1 and $1$, over $16$ dimensions. We do not need any amplification of the latent vector (i.e., we multiply by $1$, whereas the original work multiplies it by $100$). At each training step, a batch of instances is sampled from the training distribution, a batch of latent vectors is sampled from that latent space, the conditioned policy is evaluated for each instance, for each latent vector, and only the best performing latent vectors, for each instance, are used in the computation of the loss. This creates the specialisation (diversification). For additional motivation and mathematical formulation, we refer the reader to chalumeau2023combinatorial. Like all training processes in our experimental study, the method is allowed to train until convergence. In practice, each compass checkpoint was trained for $100$ million steps.
Appendix G Pseudo-algorithms
In this section, we provide algorithmic descriptions for the inference strategies used in the paper, showing how the policy is used at inference time, under the time and budget constraints. Stochastic sampling is explained in Algorithm˜1, sgbs in Algorithm˜2, online fine-tuning in Algorithm˜3 and compass in Algorithm˜4.
@@ Appendix G Pseudo-algorithms
3:while elapsed time $T_{\max}$ do
@@
3:while elapsed time $T_{\max}$ do
@@
4:while elapsed time $T_{\max}$ do
@@
4:while elapsed time $T_{\max}$ do
@@
Appendix H Real-world examples where inference-time search is applicable
Applying RL in real-world applications is a key driver of this research. Below, we describe three real-world cases where we have first-hand experience. In all of these, the user allows for a time budget, and a proxy of the solution’s quality is accessible to inform the search.
Printed Circuit Board Routing
The user submits the specification of a PCB board they want to route. This task can take up to several days or even weeks, even for an expert, depending on the complexity of the board. Therefore, users are typically satisfied with allowing an AI model to take at least a few minutes or even hours for search to obtain a better solution (routed board). Additionally, the user will often find it beneficial to leave more time and to grant more computation in order to improve the board further, which could then be easier to manufacture or reduce its production cost.
Bin Packing
When a logistics company wants to load containers with objects that need to be shipped overseas, operators typically have a few hours or days to plan which objects need to be transported, and can hence allow inference-time search. Here as well, the energy and money gains from having a better object dispatching (thus fewer containers to ship for the same amount of objects) will always justify a few hours of search, or the allowance of more computing spend.
Train Scheduling
Designing train schedules to meet a demand forecast, while proposing a feasible schedule (e.g., no collisions between trains), is another relevant example. This process usually happens once or twice a year and can take months to complete. Hence, there is no need to rely on a single rollout of the trained policy, and an inference-time search can instead be conducted over several days.
Appendix I Parameter counts for all base policy checkpoints
Creating a COMPASS checkpoint from a pre-trained base policy requires adding new parameters to process the latent vectors (see Section˜F.2), hence increasing the model size. However, this increase is small, never exceeding 2% of the total base policy model size.
We report the parameter count for all base policy checkpoints and all corresponding COMPASS checkpoints for IPPO (Table˜9), MAPPO (Table˜8) and Sable (Table˜10). We observe that (i) for IPPO and MAPPO, the increase in model size stays below 1.6%, and (ii) for Sable, it typically adds 0.5% to the total parameter count. Consequently, this model size increase has a minor impact on the memory and computational footprint of the method, which is confirmed by our study of the attempts achieved over time (reported in Appendix˜J).
Table 8: Parameter count comparison for MAPPO.
Scenario Base Policy COMPASS Policy % Increase
small-4ag-hard 125061 127109 1.64%
medium-6ag 125317 127365 1.63%
tiny-4ag-hard 125061 127109 1.64%
xlarge-4ag 125061 127109 1.64%
xlarge-4ag-hard 125061 127109 1.64%
large-8ag 125573 127621 1.63%
tiny-2ag-hard 124805 126853 1.64%
large-8ag-hard 125573 127621 1.63%
medium-4ag 125061 127109 1.64%
large-4ag 125061 127109 1.64%
large-4ag-hard 125061 127109 1.64%
medium-4ag-hard 125061 127109 1.64%
smacv2_20_units 187417 189465 1.09%
smacv2_10_units 151567 153615 1.35%
small-4ag 125061 127109 1.64%
con-10x10x10a 124293 126341 1.65%
con-15x15x23a 125957 128005 1.63%
Table 9: Parameter count comparison for IPPO.
Scenario Base Policy COMPASS Policy % Increase
large-8ag-hard 125573 127621 1.63%
medium-6ag 125317 127365 1.63%
large-8ag 125573 127621 1.63%
tiny-4ag-hard 125061 127109 1.64%
small-4ag-hard 125061 127109 1.64%
tiny-2ag-hard 124805 126853 1.64%
xlarge-4ag-hard 125061 127109 1.64%
large-4ag-hard 125061 127109 1.64%
large-4ag 125061 127109 1.64%
xlarge-4ag 125061 127109 1.64%
medium-4ag 125061 127109 1.64%
medium-4ag-hard 125061 127109 1.64%
smacv2_20_units 187417 189465 1.09%
smacv2_10_units 151567 153615 1.35%
small-4ag 125061 127109 1.64%
con-10x10x10a 124293 126341 1.65%
con-15x15x23a 125957 128005 1.63%
Table 10: Parameter count comparison for SABLE.
Scenario Base Policy COMPASS Policy % Increase
large-8ag 390096 392144 0.52%
smacv2_20_units 43363 43875 1.18%
con-10x10x10a 388422 390470 0.53%
con-15x15x23a 389907 391955 0.53%
smacv2_10_units 201435 202459 0.51%
medium-6ag 389454 391502 0.53%
medium-4ag-hard 389580 391628 0.53%
large-4ag-hard 389004 391052 0.53%
large-8ag-hard 389520 391568 0.53%
small-4ag-hard 389004 391052 0.53%
tiny-4ag-hard 1080524 1082572 0.19%
small-4ag 389196 391244 0.53%
xlarge-4ag 389004 391052 0.53%
medium-4ag 389580 391628 0.53%
xlarge-4ag-hard 1078796 1080844 0.19%
large-4ag 389004 391052 0.53%
tiny-2ag-hard 388746 390794 0.53%
Appendix J Discussion about the inference-time budget
@@ Appendix I Parameter counts for all base policy checkpoints
To remain consistent with previous literature, and provide a better understanding of the search and adaptation cost of each inference strategy and each base policy, we provide the number of attempts achieved within the time limit, for all tasks and compute capacity (i.e., number of parallel attempts allowed). Values for sable within 300 seconds are reported on Table˜11 for compass, Table˜12 for online fine-tuning, Table˜13 for stochastic sampling. sgbs builds new solutions from existing partial solutions, hence the count of the attempted solution cannot be obtained in the same way as other methods.
Values for all base policies within 30 seconds are reported on Table˜14Table˜16 and Table˜15.
@@
Appendix K Additional discussion about COMPASS
In Section˜4.3 we mention explanations to why compass gets good performance and scaling trends. In this section, we elaborate on these, and highlight three main aspects.
First, compass trains a “multiple attempts” objective, unlike other methods, which are training for “single attempt” performance. Thus, compass ’s training objective anticipates that what matters is the maximum performance over multiple attempts, rather than the average performance over these attempts. While its average solution quality may be lower, it has a higher probability of producing one excellent solution. This makes it inherently better suited for inference-time search where multiple attempts are allowed.
Second, compass creates a dense space of diverse and specialised policies, hence creating a strong exploration capacity. While most other methods rely on the entropy of the action distribution to explore (i.e. sampling stochastically from the pre-trained policy), compass can use the wide diversity encoded in its latent space to build much more diverse policies.
Finally, the Covariance Matrix Adaptation mechanism (cma-es) used by compass at inference time enables it to navigate the latent space and identify the most adapted (suitable) latent for the problem instance at hand. Using a cma-es search in a 16-dimensional space is more efficient and less prone to getting stuck in local optima compared to using gradient descent in the entire policy parameter space (i.e. online fine-tuning).
Appendix L Additional experimental details and error bars
In Fig.˜4, we report the zero-shot performance of SOTA method sable, and compare it with the best results observed for 30 seconds of inference-time search. We report the error values in Table˜17 and the figure with error bars in Fig.˜12.
We report the mean and 95% bootstrap confidence intervals. The 20M timestep values are aggregated over 10 independent runs and computed directly from the benchmarking data provided along with the Sable paper [mahjoub2025sable], for the policy trained to convergence and for COMPASS, we give the results aggregated over 128 independent runs.
Figure 12: Training the SOTA algorithm Sable to convergence is not enough to achieve optimal zero-shot performance, but using 30 seconds of search enables to consistenly reach above 85% normalised performance across the whole benchmark. We report the mean and 95% bootstrap confidence intervals.
Table 17: Performance comparison between the base policy (after 20M timesteps and after convergence) and COMPASS. Values indicate mean performance with 95% confidence intervals.
Scenario 20M Timesteps Converged COMPASS
con-10x10x10a 0.388 (0.372, 0.404) 0.500 (0.414, 0.586) 0.953 (0.914, 0.984)
con-15x15x23a 0.075 (0.067, 0.083) 0.078 (0.039, 0.125) 0.867 (0.805, 0.922)
large-4ag 0.246 (0.200, 0.283) 0.642 (0.618, 0.664) 0.949 (0.940, 0.957)
large-4ag-hard 0.175 (0.113, 0.225) 0.609 (0.581, 0.633) 0.933 (0.922, 0.943)
large-8ag 0.203 (0.197, 0.208) 0.496 (0.464, 0.526) 0.953 (0.947, 0.959)
large-8ag-hard 0.209 (0.202, 0.216) 0.516 (0.489, 0.541) 0.957 (0.950, 0.963)
medium-4ag 0.376 (0.346, 0.396) 0.666 (0.642, 0.686) 0.947 (0.941, 0.953)
medium-4ag-hard 0.253 (0.220, 0.282) 0.701 (0.677, 0.721) 0.949 (0.943, 0.955)
medium-6ag 0.240 (0.227, 0.250) 0.702 (0.668, 0.733) 0.958 (0.954, 0.963)
smacv2_10_units 0.648 (0.623, 0.675) 0.852 (0.789, 0.914) 0.977 (0.945, 1.000)
smacv2_20_units 0.599 (0.548, 0.639) 0.719 (0.641, 0.797) 0.969 (0.938, 0.992)
small-4ag 0.384 (0.243, 0.484) 0.728 (0.706, 0.746) 0.963 (0.956, 0.969)
small-4ag-hard 0.349 (0.327, 0.371) 0.707 (0.687, 0.725) 0.955 (0.948, 0.962)
tiny-2ag-hard 0.664 (0.646, 0.684) 0.810 (0.799, 0.821) 0.978 (0.967, 0.989)
tiny-4ag-hard 0.548 (0.529, 0.566) 0.783 (0.756, 0.806) 0.958 (0.951, 0.964)
xlarge-4ag 0.193 (0.118, 0.261) 0.485 (0.465, 0.504) 0.905 (0.893, 0.916)
xlarge-4ag-hard 0.065 (0.001, 0.161) 0.394 (0.373, 0.413) 0.892 (0.879, 0.906)